Q:-The displacement of two identical particles executing SHM are represented by equations x1=4sin(10t+(π/6)) and x2=5cos(ωt) . For what value of ω energy of both the particles is same ?


A. 16 units


B. 6 units


C. 4 units


D. 8 units

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Tapasranjan Das

total energy os particle executing SHM is = 1/2 mw2A2

 for 1 st perticle  E1 = 1/2 m (10)2(4)2

 for 2nd  perticle  E2 = 1/2 m (w)2(5)2 

 E1 =E2

 so w = 10*4/5

          =8 

     Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin